OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. ANSWERS BELOW PLS HELP

OpenStudy (anonymous):

y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (anonymous):

All of the equations have a slope of -3, so that part of the question is easy. The next step is to evaluate of point (7, -2) is on the line. The easy way is to fill out the x and y coordinate in each description of the line and see if it fits. A quick review of the functions shows that the third one may look promising: filling out x=7 means that the right hand side of the equation drops to 0. That implies that y+2=0, so y=-2. This is exactly the point we were looking for. Option 3 therefore.

OpenStudy (anonymous):

thanks